larry mitchel invested part of his 29,000 advance at 7% annual simple interest and the rest at %4 if his yearly interest from both is $1610 find the amount invested at each
Answer by fractalier(6550)

Call what he invested at 7%, x.
That means he invested 29000-x at 4%.
The set up is like this
.07x + .04(29000-x) = 1610
Now solve for x...
.07x + 1160 - .04x = 1610
.03x = 450
x = $15000 at 7%
29000-x = $14000 at 4%